OpenStudy (anonymous):

by 8:00am 65% of the team had reported for practice. What is the freaction of the team had not yet reported for practice?

OpenStudy (bahrom7893):

100/100-65/100 = 35/100 = .35

OpenStudy (anonymous):

100%-65%=35% as a fraction 35/100 which reduces to 7/20